Opened 7 years ago

Last modified 15 months ago

#353 assigned flight processing

EU09/02, flight day 212b/2010, Alps

Reported by: ella Owned by: knpa
Priority: alpha 4 medium Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: knpa

Description (last modified by knpa)

Data location: ~/arsf_data/2010/flight_data/archive-overflow/EU09_02-2010_212b_Alps
Project has been fast copied to ~arsf/workspace/EU09_02-2010_212b_Alps

Data arrived from ARSF via network transfer on 9th August 2010.

Scientific objectives: This project will utilise an integrated remote sensing dataset to:

  • resolve the spatial extent of glacial debris and determine debris thickness;
  • contribute to improvements in glacier mass balance modelling;
  • investigate up-scaling of field to satellite measurements, to determine the accuracy of our remote technique.

Priority: a4

PI: Dr David Rippin

Sensors:

  • Eagle (24/05/11)
  • Hawk (24/05/11)
  • Leica LIDAR (22/09/10)
  • Leica RCD (16/09/10)

Change History (43)

comment:1 Changed 7 years ago by ella

  • Description modified (diff)

comment:2 Changed 7 years ago by ella

This flight was flown on the same day as GB070/09-212a and the two study areas are less than 11 minutes fling time apart, so the Applanix and Ipas navigation were processed once and sbet/sol files copied into each project directory.

comment:3 Changed 7 years ago by jb362

The steep terrain for this project area means that an SRTM dem will be useless here and therefore hyperspectral processing should be held until a LiDAR dem has been produced.

comment:4 Changed 7 years ago by ella

Starting camera processing.

comment:5 Changed 7 years ago by jb362

Image event file has been edited for images which failed to
record a time (using time differences from sensorstat log). Images
affected are 1-5

comment:6 Changed 7 years ago by ella

Beginning camera delivery check.

comment:7 Changed 7 years ago by jb362

I actually did the camera processing for this project. Ella put a note on the ticket that she was starting the processing because I forgot to!

comment:8 Changed 7 years ago by ella

Camera check complete. Everything looks good, just need to move the camera_delivery.log file out of the delivery directory.

comment:9 Changed 7 years ago by jb362

Camera ready for delivery.

comment:10 Changed 7 years ago by jb362

Iopa will be delivering this for me since I'm leaving.

comment:11 Changed 7 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

Processing the LiDAR for this project along with Alps 212a

comment:12 Changed 7 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:13 Changed 7 years ago by emca

LiDAR data ready for delivery check at
~airborne/workspace/EU09_02-2010_212b_Alps/delivery/20100910/EU09_02

comment:14 Changed 7 years ago by iopa

Checking LiDAR delivery.

comment:15 Changed 7 years ago by iopa

LiDAR delivery check; nothing wrong to report, ready to go.

comment:16 Changed 7 years ago by emca

LiDAR data sent to PI 22/09/2010

comment:17 Changed 7 years ago by emca

  • Description modified (diff)

comment:18 Changed 7 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:19 Changed 7 years ago by benj

  • Description modified (diff)

comment:20 Changed 7 years ago by knpa

  • Other processors set to knpa

Taking HS.

comment:21 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Owner changed from emca to knpa
  • Status changed from accepted to assigned

Logsheet incorrectly gives project code as EU09_07, have corrected.

comment:22 Changed 7 years ago by knpa

SCTs

Note: There were no straight lines to dewobble in this project so these were estimated from overlaying with the LiDAR.

Eagle Hawk
1 0.02 0.03
2 0.03 0.03
3 0.04 0.04
4 0.03 0.06

comment:23 Changed 7 years ago by knpa

The Eagle and Hawk lines have large wobbles/distortions at the ends of the flightlines. It looks like they switched off after the lidar, when the plane was starting to turn.

comment:24 Changed 7 years ago by knpa

There is a serious issue of misalignment (in a direction perpendicular to flight path, mostly in specific places). Doesn't look it's due to boresight or sct but is something to do with abrupt changes in height. Investigating.

comment:25 Changed 7 years ago by knpa

The above issue was fixed by regenerating the DEM (LIDAR patched with ASTER) at 1m resolution.

comment:26 Changed 7 years ago by knpa

Hyperspectral delivery created (20110518). Ready for checking.

comment:27 Changed 7 years ago by mark1

Delivery has been checked. All data viewed and looks good.

Couple of things to check prior to delivery:

Overflows on level1 data should be mentioned in the readme
DEM name is different in directory and example command
Is the geoid-spheroid grid required? I assume the DEM is wgs84 referenced in which case it is not and should be removed from the example command.

comment:28 Changed 7 years ago by knpa

Above issues have been corrected. Preparing for dispatch.

comment:29 Changed 7 years ago by knpa

  • Description modified (diff)

Hyperspectral delivered today to David Rippin in York (24/05/11)

comment:30 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving

Project is r-synced. Ready for archiving.

comment:31 Changed 6 years ago by knpa

Started reprocessing hyperspectral with APL.

comment:32 Changed 6 years ago by knpa

Finished stage 2 - need to check lev3 output.

comment:33 Changed 6 years ago by knpa

Lev3 output looks OK.

Awaiting hawk bad pixel file.

comment:34 Changed 6 years ago by knpa

Reprocessing as apl chain has been updated to generate additional masked files.

comment:35 Changed 6 years ago by knpa

Re-ran stage 2 for Hawk as aplcal has been updated in relation to the bad pixel file.

comment:36 Changed 5 years ago by knpa

Stage 3 completed - ready for delivery check.

comment:37 Changed 5 years ago by anhi

Reprocessing delivery check

  • No Readme
  • No dem header file

comment:38 Changed 5 years ago by knpa

Located header file and generated ReadMe.

ReadMe needs to be checked.

comment:39 Changed 5 years ago by anhi

Continued reprocessing delivery check.

  • Logsheet needs renaming to EU09_02-2010_212b-Alps.pdf (_ to - in front of Alps)

Everything else looks good.

comment:40 Changed 5 years ago by knpa

Have renamed logsheet.

A re-delivery has not been requested for this flight, so this is finished with.

comment:41 Changed 4 years ago by knpa

Merged and deleted workspace version.

comment:42 Changed 18 months ago by dac

Archiving

Reprocessed hyperspectral data available from NEODC. Started uploading raw, lidar and camera.

comment:43 Changed 15 months ago by dac

Archiving

LiDAR, Photography (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EU09_02/EU09_02-2010_212b_Alps) and raw data now available from NEODC.

Note: See TracTickets for help on using tickets.